# Dark matter halo as a source of regular black-hole geometries

### v1

@@ Abstract
We construct exact black-hole solutions free of curvature singularities, sourced by dark matter halos described by galactic density profiles. Regularity of the geometry is ensured by adopting the relation $P_{r}=-\rho$ between radial pressure and density, which is consistent with the phenomenological freedom of halo models. In particular, the sufficiently dense Einasto and Dehnen-type profiles for the dark matter halo can produce asymptotically flat solutions of singularity-free black holes embedded in the galactic environment. The resulting regular black holes surrounded by dark matter are shown to be stable under axial perturbations. We further compute the shadow radii and Lyapunov exponents associated with the photon circular orbits around these black holes.
@@ I Introduction
The resolution of singularities predicted by classical general relativity remains one of the central challenges in black-hole physics. According to the singularity theorems, geodesic incompleteness is inevitable under broad physical assumptions, suggesting that classical black holes conceal regions where the known laws of physics break down. A wide variety of approaches have been developed to avoid singularities, ranging from quantum-gravity inspired modifications of Einstein’s equations to effective models in which exotic matter sources regularize the geometry Bardeen (1968); Hayward (2006); Ansoldi (2008). A key question in all such constructions is whether the required matter sources can be justified within astrophysical settings.
@@
On the other hand, black holes in nature are never isolated but always reside in astrophysical environments dominated at large scales by dark matter. Overwhelming observational evidence supports the existence of galactic and cluster dark matter halos Navarro et al. (1997); Bertone et al. (2005), although the microscopic nature of dark matter remains unknown. Phenomenological halo models, such as the Hernquist, Navarro–Frenk–White (NFW), and other profiles Hernquist (1990); Navarro et al. (1997), are widely used in galactic dynamics and cosmological simulations. Importantly, these models prescribe the density distribution without fixing a unique relation between pressure and density, thereby leaving freedom for effective fluid descriptions.
@@
In this work we exploit this freedom to construct exact black-hole solutions sourced by dark matter halos, focusing on the Einasto Einasto (1965) and Dehnen-type Dehnen (1993) density profiles. We show that assuming the radial pressure satisfies $P_{r}=-\rho$ leads to spherically symmetric spacetimes that are everywhere regular and asymptotically flat, with horizons forming under suitable parameter conditions. Because the phenomenological profiles specify only the density, our choice of pressure does not conflict with halo phenomenology, while ensuring regularity of the interior. The resulting solutions can be interpreted as singularity-free black holes embedded in galactic environments. We will also show that the obtained configuration of the regular black hole and halo is stable against axial gravitational perturbations of spacetime. As a basic geometry probe of the obtained spacetime, we calculate the shadow radii for several representative examples of regular black holes.
@@
The work is organized as follows. In Section II we formulate the general approach for constructing regular black-hole geometries sourced by spherically symmetric dark-matter distributions. Sections III and IV present explicit examples of such solutions, respectively, for the Einasto and Dehnen-type density profiles. The stability of the obtained configurations under axial perturbations is analyzed in Section V. The shadows cast by the obtained black holes and the Lypunov exponents for null circular geodesics are considered in Section VI. In Section VII we discuss the physical interpretation and astrophysical relevance of these results, while the main conclusions are summarized in Section VIII.
@@
II Spherically symmetric dark matter distributions and generic regular black-hole solutions
@@
which automatically satisfies Eq. (7). While this condition is more naturally associated with dark energy than with dark matter on cosmological scales, several studies have explored fluids obeying $p=-\rho$, or approaching this limit in certain regimes, as possible unified descriptions of dark matter and dark energy Kamenshchik et al. (2001); Bento et al. (2002); Scherrer (2004). Related concepts arise in models of vacuum-like or self-interacting dark matter, in which the dark component effectively behaves as a cosmological-constant term within overdense regions or compact objects Balakin et al. (2003); Sahni (2004). It should be emphasized, however, that in our context this equation of state is assumed to hold only within a central part of a galaxy.
@@ III Einasto profile
An example of a density distribution obeying all the above conditions is the Einasto profile Einasto:1965; Einasto and Haud (1989); Retana-Montenegro et al. (2012):
@@
Having in mind a scenario in which supermassive galactic black holes could be formed during a long period of time as a result of merging of many smaller black holes, consideration of the Einasto profile does not sound completely exotic. It is shown in Baes (2022) that only the Einasto index $n\geq 1/2$ can be supported by an isotropic orbital structure. The Einasto profile Einasto:1965; Einasto and Haud (1989) has become one of the most successful empirical descriptions of dark-matter halos, accurately reproducing the smooth curvature of the density slope seen in dynamical simulations Merritt et al. (2006); Navarro et al. (2004); Springel et al. (2008); Acharyya et al. (2024). Subsequent works established its close connection to the Sérsic law for stellar systems and related the profile’s shape parameter to the halo’s mass-accretion history Graham et al. (2006); Ludlow et al. (2013).
It is worth emphasizing that the Einasto profile defines only the density and leaves the pressure profile undetermined. In particular, there is no intrinsic equation of state associated with the Einasto model, since the relation between pressure and density depends on the chosen dynamical description of the halo (e.g., collisionless particles with velocity dispersion governed by the Jeans equations, or an effective fluid approximation). Therefore, adopting the condition $P_{r}(r)=-\rho(r)$ for the radial pressure is not in conflict with the standard assumptions underlying the Einasto halo, but rather represents a phenomenological choice that ensures regularity of the black-hole solution while remaining consistent with the absence of a fixed pressure–density relation in the halo model.
@@
Figure 1: Metric functions for the black-hole solutions with Einasto profile $n=1/2$ (left panel) and $n=1$ (right panel): $h=0.1M$ (blue), $h=0.2M$ (green), and $h=0.3M$ (red).
@@
Figure 2: Metric functions for the black-hole solutions with Einasto profile for large $n$.
@@
Using the above general procedure, we now consider several examples of density profiles that lead to regular black-hole solutions. In certain cases, these profiles admit simple analytic expressions, while in the general case the mass function must be obtained numerically by evaluating the corresponding integral. Both analytic and numerical examples are presented below. The accompanying Mathematica® notebook provides the code for generating solutions for arbitrary values of the profile parameters.
@@
$\bullet$ For other values of $n$ one can calculate the integral (11) numerically. Then, studying various numerical solutions at different values of the parameters, we observe that the black-hole solutions exist only for sufficiently dense profiles, corresponding to small values of $h$. The maximum value of $h/M$, for which the event horizon exists, decreases with $n$.
For small $n$, the halo density decreases fast and the black-hole metric outside the horizon deviates from the Schwarzschild geometry only slightly, unless $h$ is sufficiently close to its extreme value (see Fig. 1).
@@
Figure 3: Shadow radius (left panel) and Lyapunov exponents (right panel) as functions of the scale parameter $h$ in units of its maximum value $h_{m}$ for the black-hole solutions with Einasto profile (from top to bottom): $n=1$ (cyan), $n=2$ (blue), $n=3$ (green), $n=4$ (orange), $n=5$ (red), $n=6$ (magenta).
@@ IV Dehnen-type distributions
The density becomes infinite at $r=0$, unless $\alpha\leq 0$. We shall also consider $\gamma>3$, which satisfies the condition (14). Then, in order to have a monotonously decreasing profile (12), which satisfies the weak energy conditions, $\alpha=0$. In this case the total mass is
@@
Again, for sufficiently small values of $a/M$, the solution possesses an event horizon.
@@ V Axial perturbations
is a positive self-adjoint operator in the Hilbert space of square integrable functions of $r_{*}$. Therefore, all solutions of the perturbation equations with compact support initial conditions are bounded Konoplya and Zhidenko (2011), that is, the configuration is stable against axial-type gravitational perturbations.
@@ VI Black–hole shadow and Lyapunov exponent
One of the most important observable characteristics of a black hole surrounded by its galactic environment is the radius of its shadow. Consequently, a substantial body of research has been devoted to studying black hole shadows in various theories of gravity and for different types of surrounding matter and environments (see  Cunha and Herdeiro (2018); Perlick and Tsupko (2022); Bambi (2017); Hioki and Maeda (2009); Bonanno et al. (2025); Konoplya (2019); Younsi et al. (2016); Bambi et al. (2019); Konoplya et al. (2025a); Perlick et al. (2015); Konoplya and Stashko (2025); Khodadi et al. (2020); Kumar and Ghosh (2020); Konoplya and Zhidenko (2021); Afrin et al. (2021); Zakharov (2014, 2025a, 2025b); Lütfüoğlu (2025); Konoplya et al. (2020); Konoplya and Zhidenko (2019); Tsukamoto (2018) and literature therein for reviews and recent examples.
@@
Notice that $B(r)$ drops out of Eqs. (39)–(40); only $f(r)$ determines the photon sphere and the shadow radius $R_{S}$.
@@
Figure 4: Shadow radius (left panel) and Lyapunov exponents (right panel) as functions of the scale parameter $a$ for the black-hole solutions with Dehnen-type profiles for $\alpha=0$ (from shorter to longer region of $a$): $\gamma=4$, $k=1$ (cyan), $\gamma=4$, $k=2$ (blue), $\gamma=4$, $k=3$ (green), $\gamma=5$, $k=1$ (orange), $\gamma=5$, $k=2$ (red), $\gamma=5$, $k=3$ (magenta).
For Dehnen-type profiles with $\alpha=0$ (see Fig. 4), the halo density is higher for smaller values of $\gamma$ and also slightly increases as the parameter $k$ decreases. Consequently, the range of black hole shadow radii becomes larger for smaller $\gamma$ and $k$, whereas the presence of the event horizon is observed as the scalr parameter $a$ varies within a comparatively narrower interval. In contrast, the Lyapunov exponent exhibits a broader variation range for higher values of $\gamma$, indicating stronger sensitivity of the photon orbit stability to changes in the halo profile.
It should be noted that the computed shadow radii and Lyapunov exponents can also be used to estimate quasinormal modes in the eikonal (high-multipole) limit through the well-known correspondence between the orbital frequency/shadow’s radius of the circular null geodesic and the real part of the quasinormal frequency, as well as between the Lyapunov exponent and the damping rate Cardoso et al. (2009); Jusufi (2020). Although several exceptions to this correspondence have been identified in Konoplya and Stuchlík (2017); Khanna and Price (2017); Konoplya (2023); Bolokhov (2024); Konoplya and Zinhailo (2020); Konoplya et al. (2025b), the present wave equation exhibits proper WKB eikonal behavior, ensuring that the correspondence remains valid in our case.

### v3

@@ Abstract
We construct exact black hole solutions free of curvature singularities, sourced by dark matter halos described by galactic density profiles. Regularity of the geometry is ensured by adopting the relation $P_{r}=-\rho$ between radial pressure and density, which is consistent with the phenomenological freedom of halo models. Under the assumptions of regularity and the weak-energy condition, sufficiently dense dark matter halos can give rise to asymptotically flat, singularity-free black holes embedded in a galactic environment. These regular black holes are shown to be stable under axial perturbations. In particular, we obtain solutions corresponding to Einasto and Dehnen-type dark matter profiles. We further compute the shadow radii and Lyapunov exponents associated with photon circular orbits around these black holes.
@@ I Introduction
The resolution of singularities predicted by classical general relativity remains one of the central challenges in black-hole physics. According to the singularity theorems, geodesic incompleteness is inevitable under broad physical assumptions, suggesting that classical black holes conceal regions where the known laws of physics break down. A wide variety of approaches have been proposed to resolve the central singularities of classical black holes. One class of models is based on phenomenological regular metrics, where the spacetime geometry is modified at short distances in such a way that curvature invariants remain finite, as in the constructions by Bardeen, Hayward, Simpson and Visser Bardeen (1968); Hayward (2006); Simpson and Visser (2019). Another line of research employs effective matter sources, often interpreted as nonlinear electrodynamics or anisotropic fluids, which regularize the geometry while preserving asymptotic flatness Ansoldi (2008); Ayon-Beato and Garcia (1998); Dymnikova (1992); Bronnikov (2001, 2024); Bronnikov and Fabris (2006). Regular black holes have also emerged in quantum-gravity-inspired frameworks, including loop quantum gravity, asymptotically safe gravity, quasi-topological and nonlocal or higher-derivative extensions of general relativity, where quantum corrections eliminate classical singularities Kazakov and Solodukhin (1994); Modesto (2009); Bonanno and Reuter (2000); Lan et al. (2023); Bonanno et al. (2024, 2025); Spina (2025); Zhang et al. (2025); Solodukhin and Tagiev (2025); Bueno et al. (2025a, b, c); Frolov and Zelnikov (2026). Various regular black-hole solutions can be obtained in brane-world scenarios Bronnikov et al. (2007, 2003); Casadio et al. (2002). More recently, systematic studies have examined the dynamical and radiative properties of regular black holes, including accretion, grey-body factors, Hawking radiation, and, most notably, quasinormal modes, revealing both qualitative similarities and quantitative deviations from their singular counterparts Bronnikov et al. (2012); Fernando and Correa (2012); Flachi and Lemos (2013); Li et al. (2013); Liu et al. (2020); Jusufi et al. (2021); Toshmatov et al. (2015, 2018, 2019); Rincón and Santos (2020); Yang et al. (2021); Franzin et al. (2022); Konoplya et al. (2022); Meng and Zhang (2023); Franzin et al. (2024); Konoplya et al. (2023); Bolokhov (2024a); Konoplya (2025); Skvortsova (2025); Pedrotti and Vagnozzi (2024); Stashko (2024); Skvortsova (2024); Khoo (2025); Konoplya and Zhidenko (2025); Bolokhov and Skvortsova (2025a, b); Arbelaez (2025).
@@
A key question in all such constructions is whether the required matter sources can be justified within astrophysical settings. Indeed, black holes in nature are never completely isolated but always reside in astrophysical environments dominated at large scales by dark matter. Overwhelming observational evidence supports the existence of galactic and cluster dark matter halos Navarro et al. (1997); Bertone et al. (2005), although the microscopic nature of dark matter remains unknown. Phenomenological halo models, such as the Hernquist Hernquist (1990), Navarro–Frenk–White (NFW) Navarro et al. (1997), and other profiles, are widely used in galactic dynamics and cosmological simulations. Importantly, these models prescribe the density distribution without fixing a unique relation between pressure and density, thereby leaving freedom for effective fluid descriptions.
In this work we exploit this freedom to construct exact black hole solutions sourced by dark matter halos, focusing on the Einasto Einasto (1965a) and Dehnen-type Dehnen (1993) density profiles. Regular black hole configurations associated with the Dehnen-type profile, which was referred to as the Dekel-Zhao profile Dekel et al. (2017); Zhao (1996), were recently reviewed in Kar and Kar (2025), with particular emphasis on solutions with infinite asymptotic mass. The barotropic subclass of such configurations was also discussed in Sajadi et al. (2024, 2025).
@@
In the present study, we establish the general criteria and delineate the characteristic features of density profiles that give rise to regular black-hole geometries. We show that assuming the radial pressure satisfies $P_{r}=-\rho$ leads to spherically symmetric spacetimes that are everywhere regular and asymptotically flat, with horizons forming under suitable parameter conditions. Within this framework, we obtain a new family of exact solutions describing regular black holes immersed in galactic halos for a broad class of density profiles. In particular, we derive explicit and remarkably simple analytic asymptotically flat black-hole metrics for the Dehnen-type and Einasto distributions in specific parameter ranges. Since the phenomenological profiles specify only the density, our choice of pressure does not conflict with halo phenomenology, while ensuring regularity of the interior. The resulting solutions can be interpreted as singularity-free black holes embedded in galactic environments. We will also show that the obtained configuration of the regular black hole and halo is stable against axial gravitational perturbations of spacetime. As a basic geometry probe of the obtained spacetime, we calculate the shadow radii for several representative examples of regular black holes.
@@
The work is organized as follows. In Section II we formulate the general approach for constructing regular black-hole geometries sourced by spherically symmetric dark-matter distributions. Sections III and IV present explicit examples of such solutions, respectively, for the Einasto and Dehnen-type density profiles. The stability of the obtained configurations under axial perturbations is analyzed in Section V. The shadows cast by the obtained black holes and the Lyapunov exponents for null circular geodesics are considered in Section VI. In Section VII we discuss the physical interpretation and astrophysical relevance of these results, while the main conclusions are summarized in Section VIII.
@@
II Spherically symmetric dark matter distributions and generic regular black hole solutions
@@
which automatically satisfies Eq. (7). While this condition is more naturally associated with dark energy than with dark matter on cosmological scales, several studies have explored fluids obeying $p=-\rho$, or approaching this limit in certain regimes, as possible unified descriptions of dark matter and dark energy Kamenshchik et al. (2001); Bento et al. (2002); Scherrer (2004); Dymnikova and Khlopov (2015). Related concepts arise in models of vacuum-like or self-interacting dark matter, in which the dark component effectively behaves as a cosmological-constant term within overdense regions or compact objects Balakin et al. (2003); Sahni and Shtanov (2003). It should be emphasized, however, that in our context this equation of state is assumed to hold only within a central part of a galaxy.
@@ III Einasto profile
An example of a density distribution obeying all the above conditions is the Einasto profile Einasto (1965b); Einasto and Haud (1989); Retana-Montenegro et al. (2012):
@@
If we take the limit $h\to 0$ while keeping the total mass (16) fixed, the density (15) vanishes for $r>0$, and the solution reduces to the Schwarzschild metric with a singularity at $r=0$.
Having in mind a scenario in which supermassive galactic black holes could be formed during a long period of time as a result of merging of many smaller black holes, consideration of the Einasto profile does not sound completely exotic. Indeed, if some form of the Einasto profile were applicable in the early Universe, the galactic environment could be viewed as providing conditions conducive to the formation of a supermassive black hole. It is shown in Baes (2022) that only the Einasto index $n\geq 1/2$ can be supported by an isotropic orbital structure. The Einasto profile Einasto (1965b); Einasto and Haud (1989) has become one of the most successful empirical descriptions of dark-matter halos, accurately reproducing the smooth curvature of the density slope seen in dynamical simulations Merritt et al. (2006); Navarro et al. (2004); Springel et al. (2008); Acharyya et al. (2024). Subsequent works established its close connection to the Sérsic law for stellar systems and related the profile’s shape parameter to the halo’s mass-accretion history Graham et al. (2006); Ludlow et al. (2013).
It is worth emphasizing that the Einasto profile defines only the density and leaves the pressure profile undetermined. In particular, there is no intrinsic equation of state associated with the Einasto model, since the relation between pressure and density depends on the chosen dynamical description of the halo (e.g., collisionless particles with velocity dispersion governed by the Jeans equations, or an effective fluid approximation). Therefore, adopting the condition $P_{r}(r)=-\rho(r)$ for the radial pressure is not in conflict with the standard assumptions underlying the Einasto halo, but rather represents a phenomenological choice that ensures regularity of the black hole solution while remaining consistent with the absence of a fixed pressure–density relation in the halo model.
@@
Figure 1: Metric functions for the black hole solutions with Einasto profile $n=1/2$ (left panel) and $n=1$ (right panel): $h=0.1M$ (blue), $h=0.2M$ (green), and $h=0.3M$ (red).
@@
Figure 2: Metric functions for the black hole solutions with Einasto profile for large $n$.
@@
Using the above general procedure, we now consider several examples of density profiles that lead to regular black hole solutions. In certain cases, these profiles admit simple analytic expressions, while in the general case the mass function must be obtained numerically by evaluating the corresponding integral. Both analytic and numerical examples are presented below. The accompanying Mathematica® notebook [^1] The Mathematica® notebook is publicly available from https://arxiv.org/src/2511.03066v1/anc/BHEinasto.nb. provides the code for generating solutions for arbitrary values of the profile parameters.
@@
$\bullet$ For other values of $n$ one can calculate the integral (11) numerically. Then, studying various numerical solutions at different values of the parameters, we observe that the black hole solutions exist only for sufficiently dense profiles, corresponding to small values of $h$. The maximum value of $h/M$, for which the event horizon exists, decreases with $n$.
For small $n$, the halo density decreases fast and the black-hole metric outside the horizon deviates from the Schwarzschild geometry only slightly, unless $h$ is sufficiently close to its extreme value (see Fig. 1). Similar black-hole geometries have been recently obtained by introducing Gaussian nonlocality in the radial coordinate of the Schwarzschild metric Boos (2025).
For large $n$ the Einasto density profile must be very dense and decays relatively slowly. Therefore, the black-hole metric is significantly different from the Schwarzschild geometry within the parametric range of $h$ providing the existence of the event horizon (see Fig. 2). As a result for larger $n$ the shadow radius and Lyapunov exponents generally differ more from their Schwarzschild values (see Fig. 3), which correspond to the limit $h\to 0$.
@@
Figure 3: Shadow radius (left panel) and Lyapunov exponents (right panel) as functions of the scale parameter $h$ in units of its maximum value $h_{m}$ for the black hole solutions with Einasto profile (from top to bottom): $n=1$ (cyan), $n=2$ (blue), $n=3$ (green), $n=4$ (orange), $n=5$ (red), $n=6$ (magenta). The value $h=0$ corresponds to the Schwarzschild black hole.
@@ IV Dehnen-type distributions
The density becomes infinite at $r=0$, unless $\alpha\leq 0$. This requirement rules out a number of widely used phenomenological halo profiles, including the NFW and Hernquist models, which correspond to $\alpha=1$, as well as the Moore profile with $\alpha=7/5$ Moore et al. (1998). However, those profiles are intended to describe the dark-matter distribution only on galactic scales. The formal divergence at the center indicates that they cannot be extrapolated down to small radii: in any physically realistic setting, the density must be regularized or modified near the origin.
We shall also consider $\gamma>3$, which satisfies the condition (14). Then, in order to have a monotonously decreasing profile (12), which satisfies the weak energy conditions, $\alpha=0$. It is interesting to note that, in this case, the fluid is barotropic Sajadi et al. (2025). The total mass is
@@
Again, for sufficiently small values of $a/M$, the solution possesses an event horizon. In the limit $a\to 0$ while keeping the total mass $M$ in Eq. (24) fixed, the solution reduces to the Schwarzschild geometry.
It is noteworthy that, although for $\gamma=3$ the total mass (24) diverges, the corresponding solutions remain asymptotically flat, whereas for $\gamma=2$ they are asymptotically characterized by a solid angle defect (see Kar and Kar (2025) for a review).
@@
For $\gamma=4$ and $k=2$, we have a regular black hole solution, which can be obtained within nonlinear electrodynamics minimally coupled to gravity Dymnikova (2004).
@@
For $\gamma=4$ we obtain the solution (3.18) of Kar and Kar (2025). For $\gamma=6$ the solution (28) coincides with the Hayward black hole Hayward (2006).
@@ V Axial perturbations
is a positive self-adjoint operator in the Hilbert space of square integrable functions of $r_{*}$. Therefore, all solutions of the perturbation equations with compact support initial conditions are bounded Konoplya and Zhidenko (2011), that is, the configuration is stable against axial-type gravitational perturbations. In other words, any such perturbation of the black hole surrounded by the fluid decays and eventually settles back to the unperturbed background configuration. Hence, no growing mode must be present in the axial sector.
@@ VI Black–hole shadow and Lyapunov exponent
One of the most important observable characteristics of a black hole surrounded by its galactic environment is the radius of its shadow. Consequently, a substantial body of research has been devoted to studying black hole shadows in various theories of gravity and for different types of surrounding matter and environments (see  Cunha and Herdeiro (2018); Perlick and Tsupko (2022); Bambi (2017); Hioki and Maeda (2009); Bonanno et al. (2025); Konoplya (2019); Younsi et al. (2016); Bambi et al. (2019); Konoplya et al. (2025a); Perlick et al. (2015); Konoplya and Stashko (2025); Khodadi et al. (2020); Kumar and Ghosh (2020); Konoplya and Zhidenko (2021); Afrin et al. (2021); Zakharov (2014, 2025a, 2025b); Lütfüoğlu (2025a); Konoplya et al. (2020); Konoplya and Zhidenko (2019); Tsukamoto (2018); Stashko (2024); Spina (2025) and literature therein for reviews and recent examples.
@@
Notice that $B(r)$ drops out of Eqs. (39) and (40); only $f(r)$ determines the photon sphere and the shadow radius $R_{S}$.
@@
Figure 4: Shadow radius (left panel) and Lyapunov exponents (right panel) as functions of the scale parameter $a$ for the black hole solutions with Dehnen-type profiles for $\alpha=0$ (from shorter to longer region of $a$): $\gamma=4$, $k=1$ (cyan), $\gamma=4$, $k=2$ (blue), $\gamma=4$, $k=3$ (green), $\gamma=5$, $k=1$ (orange), $\gamma=5$, $k=2$ (red), $\gamma=5$, $k=3$ (magenta). The value $a=0$ corresponds to the Schwarzschild black hole.
For Dehnen-type profiles with $\alpha=0$ (see Fig. 4), the halo density is higher for smaller values of $\gamma$ and also slightly increases as the parameter $k$ decreases. Consequently, the range of black hole shadow radii becomes larger for smaller $\gamma$ and $k$, whereas the presence of the event horizon is observed as the scale parameter $a$ varies within a comparatively narrower interval. In contrast, the Lyapunov exponent exhibits a broader variation range for higher values of $\gamma$, indicating stronger sensitivity of the photon orbit stability to changes in the halo profile.
It should be noted that the computed shadow radii and Lyapunov exponents can also be used to estimate quasinormal modes in the eikonal (high-multipole) limit through the well-known correspondence between the orbital frequency/shadow’s radius of the circular null geodesic and the real part of the quasinormal frequency, as well as between the Lyapunov exponent and the damping rate Cardoso et al. (2009); Jusufi (2020). Although several exceptions to this correspondence have been identified in Konoplya and Stuchlík (2017); Khanna and Price (2017); Konoplya (2023); Bolokhov (2024b); Konoplya and Zinhailo (2020); Konoplya et al. (2025b), the present wave equation exhibits proper WKB eikonal behavior, ensuring that the correspondence remains valid in our case.